# To See a World in a Spark of Neuron: Disentangling Multi-task Interference for Training-free Model Merging

### v2

@@ Appendix D Implementation Details
We executed all our experiments on Nvidia A6000 GPUs equipped with 48GB RAM. The merging experiments demonstrated highly computational efficiency, with evaluation times under 2 minutes for T5-Base, T5-Large, ViT-B/32, and ViT-L/14 models. For large language model, specifically LLaMA2, the validation process across three datasets required approximately 40 minutes per complete evaluation cycle.
@@ Appendix E Dataset Details
We utilized several datasets, and each of them comes with specific licenses. The following datasets are available under the Creative Commons License: WiC (Pilehvar and Camacho-Collados, 2019), WSC (Levesque et al., 2012), Story Cloze (Sharma et al., 2018), QuaRTz (Tafjord et al., 2019), Cars (Krause et al., 2013), and GTSRB (Stallkamp et al., 2011). Winogrande (Sakaguchi et al., 2021) and QASC (Khot et al., 2019) are distributed under the Apache License, while COPA (Gordon et al., 2012) is covered by the BSD-2 Clause License. WikiQA (Yang et al., 2015) is governed by the Microsoft Research Data License Agreement. Cosmos QA (Huang et al., 2019) is licensed under the CC BY 4.0. QuAIL (Rogers et al., 2020) and CMMLU (Li et al., 2024) are licensed under the CC BY-NC-SA 4.0. H-SWAG (Zellers et al., 2019), GSM8K (Cobbe et al., 2021), HumanEval (Chen et al., 2021), and EuroSAT (Helber et al., 2019) fall under the MIT License, and MNIST (LeCun, 1998) is licensed under the GNU General Public License.
@@
Table 12: Test set performance when merging T5-Base models on seven NLP tasks. Please refer to Section 5.1 for experimental details.
@@
Table 13: Test set performance when merging ViT-B/32 models on eight vision tasks. Please refer to Section 5.4 for experimental details.
@@
Table 14: Test set performance when merging ViT-L/14 models on eight vision tasks. Please refer to Section 5.4 for experimental details.
@@
Table 15: Out-of-Distribution performance of T5-Base models checkpoints on six tasks. Please refer to Section 5.2 for experimental details.
@@
Table 16: Out-of-Distribution performance of T5-Large models checkpoints on six tasks. Please refer to Section 5.2 for experimental details.
@@
Figure 8: Comparison of merging methods on NLP with T5-Large (Left) and CV with ViT-L/14 (Right) without validation datasets. NeuroMerging outperforms existing methods in most tasks. Please refer to Section 6.1 for more discussion.

### v4

@@ Appendix D Implementation Details
We executed NLP and CV experiments on Nvidia GeForce 4090 GPUs with 24GB RAM, and LLM experiments on Nvidia A6000 GPUs with 48GB RAM, respectively. The merging experiments demonstrated highly computational efficiency, with evaluation times under 2 minutes for T5-Base, T5-Large, ViT-B/32, and ViT-L/14 models. For large language model, specifically LLaMA2, the validation process across three datasets required approximately 40 minutes per complete evaluation cycle.
@@ Appendix E Dataset Details
We utilized several datasets, and each of them comes with specific licenses. The following datasets are available under the Creative Commons License: WiC (Pilehvar and Camacho-Collados, 2019), WSC (Levesque et al., 2012), Story Cloze (Sharma et al., 2018), QuaRTz (Tafjord et al., 2019), Cars (Krause et al., 2013), and GTSRB (Stallkamp et al., 2011). Winogrande (Sakaguchi et al., 2020) and QASC (Khot et al., 2020) are distributed under the Apache License, while COPA (Gordon et al., 2012) is covered by the BSD-2 Clause License. WikiQA (Yang et al., 2015) is governed by the Microsoft Research Data License Agreement. Cosmos QA (Huang et al., 2019) is licensed under the CC BY 4.0. QuAIL (Rogers et al., 2020) and CMMLU (Li et al., 2024) are licensed under the CC BY-NC-SA 4.0. H-SWAG (Zellers et al., 2019), GSM8K (Cobbe et al., 2021), HumanEval (Chen et al., 2021), and EuroSAT (Helber et al., 2018) fall under the MIT License, and MNIST (LeCun, 1998) is licensed under the GNU General Public License.
@@
Table 13: Test set performance when merging T5-Base models on seven NLP tasks. Please refer to Section 5.1 for experimental details and Section C.4 for error analysis.
@@
Table 14: Test set performance when merging ViT-B/32 models on eight vision tasks. Please refer to Section 5.4 for experimental details and Section C.4 for error analysis.
@@
Table 15: Test set performance when merging ViT-L/14 models on eight vision tasks. Please refer to Section 5.4 for experimental details and Section C.4 for error analysis.
@@
Table 16: Out-of-Distribution performance of T5-Base models checkpoints on six tasks. Please refer to Section 5.2 for experimental details and Section C.4 for error analysis.
@@
Table 17: Out-of-Distribution performance of T5-Large models checkpoints on six tasks. Please refer to Section 5.2 for experimental details and Section C.4 for error analysis.
@@
Figure 8: Comparison of merging methods on NLP with T5-Large (Left) and CV with ViT-L/14 (Right) without validation datasets. NeuroMerging outperformed existing methods in most tasks. Please refer to Section 6.1 for more discussion.